OpenStudy (anonymous):

In the following exercises, let y = safe load in pounds and x = depth in inches for a certain type of rectangular horizontal beam. A constant of proportionality k exists such that y = kx2 (y varies directly as x2). Determine the constant k for a beam with y = 1,000 pounds and x = 5 inches. K=____ What is a safe load if x = 10 inches? ____Pounds What depth should be used to support 16,000 pounds? ____inches

OpenStudy (anonymous):

y = kx^2 Putting y = 1000, x = 5 1000 = k * 5^2 1000 = 25k k = 1000 / 25 k = 40 lb / in^2 Putting x = 10 in, k = 40 lb / in^2 y = 40 * 10^2 y = 40 * 100 y = 4000 lb Putting y = 16000 lb, k = 40 lb / in^2: 16000 = 40x^2 x^2 = 16000 / 40 x^2 = 400 x = 20 in

OpenStudy (anonymous):

thanks annas :)

OpenStudy (anonymous):

np :) thats what i do...